A state highway patrol car radar unit uses a frequency of 9.50×109 Hz What frequency difference will the unit detect from a car approaching a parked patrol car with a speed of 39.0 m/s ?
Answer #1

Freq(Observed) = (V/(V-Vs) ) *Freq(source)

=C/C-39 *9.50×10^9 where C= 3*10^8m/s

=9500001235

change =1235.00015831 Hz